OpenStudy (anonymous):

cos^2 2x−cos^2 6x=sin4x sin8x please help me prove it

OpenStudy (solomonzelman):

you can manipulate both sides, right ?

OpenStudy (anonymous):

yes..but it would be prefered if we can begin with the left hand side

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \cos^2 2x−\cos^2 6x=\sin^4x \ sin^8x }\) like this ?

OpenStudy (anonymous):

\[\cos ^{2} 2x - \cos ^{2} 6x = \sin4x \sin 8x \] like this..

OpenStudy (solomonzelman):

Oh!! you know how to do !!! You figured out fast:)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \cos^2 2x−\cos^2 6x=\sin(4x) \sin(8x) }\)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \cos^2 2x−\cos^2 6x=\sin 4x \sin 8x }\) \(\normalsize\color{blue}{ \cos^2 2x−\cos^2 (4x+2x)=\sin 4x \sin 8x }\) \(\normalsize\color{blue}{ \cos^2 2x−(\cos^2 4x~~\sin^2 2x-\cos^2 2x~~\sin^2 4x)=\sin 4x \sin 8x }\) \(\normalsize\color{blue}{ \cos^2 2x−\cos^2 4x~~\sin^2 2x+\cos^2 2x~~\sin^2 4x=\sin 4x \sin 8x }\) \(\normalsize\color{blue}{ \cos^2 2x−(\cos^2 4x)~~\sin^2 2x+\cos^2 2x~~(\sin^2 4x)=\sin 4x \sin 8x }\) \(\normalsize\color{blue}{ \cos^2 2x−(\cos^2 2x-\sin^2 2x)~~\sin^2 2x+\cos^2 2x~~(2\sin 2x)=\sin 4x \sin 8x }\) \(\normalsize\color{blue}{ \cos^2 2x−\cos^2 2x+\sin^2 2x~~\sin^2 2x+\cos^2 2x~~(2\sin 2x)=\sin 4x \sin 8x }\) read this for now

OpenStudy (solomonzelman):

ignore the last step I posted

OpenStudy (anonymous):

Can I say something?

OpenStudy (solomonzelman):

No I am lost... idk, sorry

OpenStudy (solomonzelman):

OOPS go ahead

OpenStudy (anonymous):

from the left hand side: \[cos^2(2x) -cos^2(6x) = (cos (2x) -cos (6x))*(cos(2x)+cos(6x))\] so far you have 2 terms ,right?

OpenStudy (anonymous):

how is \[\cos ^{2} 4x \] = \[\cos ^{2} 2x - \sin ^{2} 2x\] ??

OpenStudy (anonymous):

I am sorry, the second term is \[cos(2x) +cos (6x)= 2 cos(\dfrac{2x+6x}{2})cos(\dfrac{2x-6x}{2})= 2 cos (4x)cos(-2x)\]

OpenStudy (anonymous):

ok..continue

OpenStudy (anonymous):

the first term is \[cos(2x) -cos (6x)= -2 sin (\dfrac{2x+6x}{2})sin(\dfrac{2x-6x}{2})= -2 sin (4x)sin(-2x)\]

OpenStudy (anonymous):

Now, times them together, because originally they are the factor of the product above, right? so at the end, you have the left hand side = 2cos(4x)cos(-2x)*(-2) sin(4x)sin (-2x)

OpenStudy (anonymous):

Now I re-arrange it as 2cos (4x) sin(4x) *(-2) cos (-2x) sin(-2x) ok so far?

OpenStudy (anonymous):

you can see \(2cos(4x) sin(4x) = sin (8x)\)

OpenStudy (anonymous):

and \(2 cos (-2x) sin(-2x) = sin (-4x) = -sin (4x)\) but in the front we have -2, not 2, so that it becomes sin(4x) --> it turns to sin(8x)sin(4x) = the right hand side

OpenStudy (anonymous):

hehehe... hopefully you can follow me. trigs is .... not hard but.... tedious!!

OpenStudy (anonymous):

thank you so much!!